OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192

OpenStudy (solomonzelman):

\(\Large\color{black}{ \displaystyle a_1=8 }\) \(\Large\color{black}{ \displaystyle a_6=-8,192 }\) for any geometric sequence, \(\Large\color{black}{ \displaystyle a_n=(a_1)\times(r)^{n-1} }\) (use this)

OpenStudy (solomonzelman):

use this to find the common ratio

OpenStudy (anonymous):

where would you put the -8192?

OpenStudy (solomonzelman):

such as, (in your case) \(\Large\color{black}{ \displaystyle a_6=(a_1)\times(r)^{6-1} }\) \(\Large\color{black}{ \displaystyle a_6=(a_1)\times(r)^{5} }\) then plug in your terms,

OpenStudy (solomonzelman):

plug in -8,192 for \(\Large\color{black}{ \displaystyle a_6}\) 8 for \(\Large\color{black}{ \displaystyle a_1 }\)